# Deep Neural Networks and Brain Alignment: Brain Encoding and Decoding (Survey)

### v2

@@ 6.4.3 NLP Tasks and Linguistic Properties in LMs and Brains
Figure 16: Comparison of brain recordings with language models trained on web corpora (Left) and language models trained on book stories (Right) (Aw & Toneva, 2023)This Figure is redrawn from Aw & Toneva (2023).
To investigate whether large language models with longer context are learning a deeper understanding of the text, Aw & Toneva (2023) used four pretrained large language models (BART, Longformer Encoder Decoder, BigBird, and LongT5) and also trained them to improve their narrative understanding, using the method detailed in Figure 16. They find that the improvements in brain alignment are larger for character names than for other discourse features, which indicates that these models are learning important narrative elements. However, it is not understood whether language models with the prediction of the next word are necessary for the observed brain alignment or simply sufficient, and whether there are other shared mechanisms or information that is similarly important. Merlin & Toneva (2022) proposed two perturbations to pretrained language models that, when used together, can control for the effects of next word prediction and word-level semantics on the alignment with brain recordings. Specifically, they found that improvements in alignment with brain recordings in two language processing regions–Inferior Frontal Gyrus (IFG) and Angular Gyrus (AG)–are due to next word prediction and word-level semantics. However, what linguistic information actually underlies the observed alignment between brains and language models was not clear. Recently, Oota et al. (2024b) tested the effect of a range of linguistic properties (surface, syntactic and semantic) and found that the elimination of each linguistic property results in a significant decrease in brain alignment across all layers of BERT. Further, syntactic properties are more responsible and have the largest effect on the trend of brain alignment across model layers. To further understand what aspects of linguistic stimuli contribute to ANN-to-brain similarity, Kauf et al. (2024b) systematically manipulated the stimuli (i.e., perturbed sentences’ word order, removed different subsets of words, or replaced sentences with other sentences of varying semantic similarity) and found that lexical semantic content rather than the sentence’s syntactic form is primarily responsible for the DNN-to-brain similarity. Similar to studies on pretrained models and brain similarity, AlKhamissi et al. (2024) investigated the reasons for the similarity of untrained language models and brain alignment by performing mechanistic interpretability of the models. By isolating components of the Transformer architecture (GPT-2 XL), they found that tokenization strategy and multihead attention are the two major components driving this better brain alignment.
@@ 6.4.4 Key Takeaways
      Contrary to previous claims in , syntax and semantics are not associated with separate modules but instead share common brain language regions and are distributed across the language network.
@@ 6.5 Auditory Encoding
Figure 17: Brain prediction using self-supervised speech model: Data2Vec. The plot shows that speech-based models better predict early auditory cortex (Oota et al., 2023c).
@@ 6.5.1 Alignment Between Pretrained Speech Models and Brains
Recently, several researchers have used popular deep learning models such as APC (Chung et al., 2020), Wav2Vec2.0 (Baevski et al., 2020), HuBERT (Hsu et al., 2021), and Data2Vec (Baevski et al., 2022) for encoding speech stimuli.  Millet et al. (2022) used a self-supervised learning model, Wav2Vec2.0, to learn latent representations of the speech waveform similar to human brain. They find that the functional hierarchy of its transformer layers aligns with the cortical hierarchy of speech in the brain and reveals the whole-brain organisation of speech processing with unprecedented clarity. This means that the first transformer layers map onto the low-level auditory cortices (A1 and A2), the deeper layers map onto brain regions associated with higher-level processes (e.g. STS and IFG). Vaidya et al. (2022) present the first systematic study to bridge the gap between recent four self-supervised speech representation methods (APC, Wav2Vec, Wav2Vec2.0, and HuBERT) and computational models of the human auditory system. Similar to (Millet et al., 2022), they find that self-supervised speech models are the best models of auditory areas. Lower layers best modeled low-level areas, and upper-middle layers were most predictive of phonetic and semantic areas, while layer representations follow the accepted hierarchy of speech processing. Tuckute et al. (2023) analyzed 19 different speech models and found that some audio models derived in engineering contexts (model applications ranging from speech recognition and speech enhancement to audio captioning and audio source separation) produce poor predictions of auditory cortical responses, many task-optimized audio speech deep learning models outpredict a standard spectrotemporal model of the auditory cortex and exhibit hierarchical layer-region correspondence with auditory cortex. Further, Oota et al. (2023c) extended this analysis to more such deep learning based speech models (30 self-supervised speech models). They found that both language and auditory brain areas, are best aligned with intermediate layers in deep learning models. As shown in Figure 17, they also found that speech models better predict early auditory cortex than late language regions. Although pretrained speech models can understand broad aspects of speech in general, the implications of finetuning speech pretrained models for various speech-processing tasks for speech encoding in the brain, remain underexplored.
Figure 18: The pretrained Wav2Vec2.0 model and finetuned to eight different downstream speech tasks and their brain alignment (Oota et al., 2023d).
@@ 6.5.2 Underlying Speech Properties in Speech Models and Brains
Understanding the reasons behind the observed similarities between speech processing in speech models and brains can lead to more insights into both systems. Recent work Oota et al. (2023d) has found that using a finetuned Wav2Vec2.0 leads to improved brain alignment. In particular, as shown in Figure 18Oota et al. (2023d) build neural speech taskonomy models for brain encoding and aim to find speech-processing tasks that have the most explanatory capability of brain activation during naturalistic story listening experiments. They find that task-specific (Automated Speech Recognition (ASR), Entity Recognition (ER), Speaker Identification (SID) and Intent Classification (IC)) speech representations lead to a significant improvement in brain alignment compared to the pretrained Wav2Vec2.0 model for specific brain regions. Finetuning on ER, SID and IC leads to the best alignment for the early auditory cortex; finetuning on ASR provides the best encoding for the auditory associative cortex and language regions. Further, the layer-wise analysis of the effect of each speech task on the alignment with whole brain activity shows that the ASR task is better aligned in middle layers. Similar to language model fine-tuning with brain data in Schwartz et al. (2019)Li et al. (2024) fine-tuned a pretrained Wav2Vec2.0 model with brain recordings and and found that this induced process modified the language representations, improving the model’s performance on downstream tasks from the SUPERB benchmark.
@@ 6.6.1 Alignment Between Vision Models and Brains
Similar to language, in vision, early models focused on independent models of visual processing (object classification) using CNNs (Yamins et al., 2014).  Eickenberg et al. (2017) use CNNs as candidate models to model human brain activity during the viewing of natural images by constructing predictive models based on their different CNN layers and BOLD fMRI activations. They find that there are similarities between the computations of convolutional networks and cognitive vision at the beginning and at the end of the ventral stream object-recognition process.  Cichy et al. (2016) further investigate the stages of human visual processing in both time (MEG recordings) and space (fMRI recordings). By comparing these findings with representations derived from deep neural networks (DNNs), the authors demonstrate that DNNs effectively encapsulate the sequential stages of human visual processing. This encompasses the progression from early visual areas towards the specialized pathways of the dorsal and ventral streams, highlighting the DNN’s capacity to mirror complex neural processes in both time and space. Despite the effectiveness of CNNs, it is difficult to draw specific inferences about neural information processing using CNN-derived representations from a generic object-classification CNN. Hence, Wang et al. (2019) built encoding models with individual feature spaces obtained from 21 computer vision tasks. One of the main findings is that features from 3D tasks, compared to those from 2D tasks, predict a distinct part of visual cortex. Recent efforts in visual encoding models, particularly self-supervised models (instance-prototype contrastive learning), operate by taking multiple samples over an image and projecting these through a deep convolutional neural network into a low-dimensional embeddings space (Konkle & Alvarez, 2022). The results show that these self-supervised models achieve parity with the category-supervised models in accounting for the structure of brain responses. Since the human visual system uses two parallel pathways for spatial processing and object recognition, while computer vision systems (CNNs) typically use a single pathway,  Choi et al. (2024) developed a dual-stream vision model to mimic human vision. This model uses two branches of CNNs to replicate the dorsal and ventral cortical pathways, aligning with the brain’s pathways and suggesting that distinct responses are driven more by visual attention and object recognition goals than by retinal input selectivity.
@@ 6.7 Multimodal Brain Encoding
Single modality stimulus. Here, participants engage in single modality stimuli, such as watching images or silent videos. Many brain encoding studies have focused on single modality stimuli, while representations are extracted from multimodal models (Oota et al., 2022e; Wang et al., 2022; Tang et al., 2024).  Oota et al. (2022e) experimented with multimodal models like CLIP, LXMERT, and VisualBERT and found VisualBERT better predict neural responses than vision-only models such as CNNs and Image Transformers. Similarly,  (Wang et al., 2022) find that multimodal models like CLIP better predict neural responses in the visual cortex than previous vision-only models like CNNs. This is attributed to the fact that high-level human visual representations encompass semantics and the relational structure of the visual world beyond object identity (Gauthier et al., 2003). Recently, Tang et al. (2024) investigated a multimodal Transformer as the encoder architecture to extract the aligned concept representations for narrative stories and movies to model fMRI responses to naturalistic stories and movies, respectively. Since language and vision rely on similar concept representations, the authors perform a cross-modal experiment in which how well the language encoding models can predict movie-fMRI responses from narrative story features (story → movie) and how well the vision encoding models can predict narrative story-fMRI responses from movie features (movie → story). Overall, the authors find that cross-modality performance was higher for features extracted from multimodal transformers than for linearly aligned features extracted from unimodal transformers.
@@
Multimodality stimulus. Here, participants engage with multi-modal stimuli (e.g., watching movies that include audio). Recent studies have built encoding models where multi-modal stimulus representations are extracted using Transformer-based multi-modal models (Dong & Toneva, 2023; Nakagi et al., 2024).  Dong & Toneva (2023) present a systematic approach to probe multimodal video Transformer model by leveraging neuro-scientific evidence of multimodal information processing in the brain. The authors find that intermediate layers of a multimodal video transformer are better at predicting multimodal brain activity than other layers, indicating that the intermediate layers encode the most brain-related properties of the video stimuli. A recent study by (Nakagi et al., 2024), which used fMRI during the viewing of 8.3 hours of video content, and discovered distinct brain regions associated with different semantic levels, highlighting the significance of modeling various levels of semantic content simultaneously. The video material was meticulously annotated in five distinct semantic categories—speech, object, story, summary, and time/place—employing advanced large language models to derive latent representations. These representations were then used to predict fMRI brain activity across the various semantic categories. The authors discovered that the lack of unique variance for Summary and TimePlace is a notable insight, suggesting that merely incorporating these types of information into encoding analyses may not adequately capture higher-level semantic representations in the brain.
@@ 7.3 Decoder Architectures
Figure 19: Scheme for Brain Decoding. Left: Image decoder (Smith, 2013), Right: Language Decoder (Wang et al., 2019).  The left Figure is adapted from Smith (2013) and the right Figure is adapted from Wang et al. (2019).
@@ 7.4 Impact of Large Models on Brain Decoding
Masked autoencoders (He et al., 2022) play a vital role in learning low-rank representations by reconstructing missing parts of the input data. In the context of brain decoding, these models are often used for pretraining by masking out some brain voxels and attempting to reconstruct them, thereby learning the underlying representations (Chen et al., 2023; 2024; Sun et al., 2024). These fMRI representations are then utilized as conditions for downstream conditional generation models, enhancing their ability to produce detailed and accurate reconstructions compared with linear models.
Contrastive learning (Khosla et al., 2020) has emerged as a powerful technique for representation learning by maximizing the similarity between related data points while minimizing the similarity between unrelated ones. This approach has been instrumental in aligning brain activity with corresponding stimuli in a shared embedding space, facilitating more accurate and semantically meaningful decoding. One of the most notable applications of contrastive learning in brain decoding is the CLIP model (Radford et al., 2021). CLIP aligns text and images in a shared embedding space, greatly enhancing the decoding of visual stimuli. These models decode brain activity into text descriptions that are then used to generate corresponding images, effectively bridging the gap between linguistic and visual representations. In brain decoding, researchers often align fMRI embeddings with CLIP-based embedding spaces, allowing for more precise and semantically rich reconstructions of visual stimuli from brain data Chen et al. (2024); Scotti et al. (2024).
@@
LLMs, particularly models in the GPT series (Brown et al., 2020), have revolutionized language decoding. These models are capable of generating coherent and contextually appropriate text based on brain activity patterns. For instance, instead of merely decoding vector representations of stimuli, recent studies have leveraged LLMs to reconstruct entire sentences or continuous language from fMRI data Tang et al. (2023); Zhao et al. (2024). This shift from vector-based decoding to full text generation has significantly enhanced the semantic richness and contextual accuracy of the decoded output. The ability of LLMs to model complex language structures and generate text conditioned on neural data has opened new avenues for understanding how the brain processes language, providing practical applications in areas such as communication aids for individuals with speech impairments.
@@
Diffusion models (Ho et al., 2020), particularly those like Stable Diffusion, have been pivotal in generating high-fidelity images from brain activity. These models leverage the noise-to-signal transformation process to produce detailed and semantically rich visual outputs. By conditioning these models on brain activity data, researchers have achieved remarkable success in reconstructing images that closely resemble the original stimuli (Scotti et al., 2024; Takagi & Nishimoto, 2022; Chen et al., 2023; 2024; Ozcelik & VanRullen, 2023; Takagi & Nishimoto, 2023). The high resolution and fidelity of the generated images represent a significant improvement over previous methods, which often struggled to capture fine details and semantic accuracy simultaneously.
@@
Continuous Language, level2, l sep=10mm,[ (Affolter et al., 2020; Tang et al., 2023), tnode] ] ] [Auditory, level1, l sep=10mm, [Speech
@@
Reconstruction , level2, l sep=10mm,[ (Naselaris et al., 2009; Beliy et al., 2019; Takagi & Nishimoto, 2022; Ozcelik & VanRullen, 2023; Chen et al., 2023; Koide-Majima et al., 2024; Takagi & Nishimoto, 2023), tnode] ] [Video
Reconstruction, level2, l sep=10mm,[ (Nishimoto et al., 2011; Le et al., 2021; Kupershmidt et al., 2022; Chen et al., 2024), tnode] ] ] ]
Figure 20: Categorization of Brain Decoding Studies
Stimuli Authors Dataset Type Lang. Stimulus Representations ||||S|||| Dataset
@@
(Chen et al., 2024) fMRI Latent Diffusion Model, CLIP 3 HCP fMRI-Video-Dataset

### v3

@@ 6.4.3 NLP Tasks and Linguistic Properties in LMs and Brains
Figure 15: Comparison of brain recordings with language models trained on web corpora (Left) and language models trained on book stories (Right). This Figure is adapted from (Aw & Toneva, 2023), and used with permission from the respective authors.
To investigate whether large language models with longer context are learning a deeper understanding of the text, Aw & Toneva (2023) used four pretrained large language models (BART (Lewis et al., 2020), Longformer Encoder Decoder (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), and LongT5 (Guo et al., 2022)) and also trained them to improve their narrative understanding, using the method detailed in Figure 15. They find that the improvements in brain alignment are larger for character names than for other discourse features, which indicates that these models are learning important narrative elements. However, it is not understood whether language models with the prediction of the next word are necessary for the observed brain alignment or simply sufficient, and whether there are other shared mechanisms or information that is similarly important. Merlin & Toneva (2022) proposed two perturbations to pretrained language models that, when used together, can control for the effects of next word prediction and word-level semantics on the alignment with brain recordings. Specifically, they found that improvements in alignment with brain recordings in two language processing regions–Inferior Frontal Gyrus (IFG) and Angular Gyrus (AG)–are due to next word prediction and word-level semantics. However, what linguistic information actually underlies the observed alignment between brains and language models was not clear. Recently, Oota et al. (2024b) tested the effect of a range of linguistic properties (surface, syntactic and semantic) and found that the elimination of each linguistic property results in a significant decrease in brain alignment across all layers of BERT. Further, syntactic properties are more responsible and have the largest effect on the trend of brain alignment across model layers. To further understand what aspects of linguistic stimuli contribute to ANN-to-brain similarity, Kauf et al. (2024b) systematically manipulated the stimuli (i.e., perturbed sentences’ word order, removed different subsets of words, or replaced sentences with other sentences of varying semantic similarity) and found that lexical semantic content rather than the sentence’s syntactic form is primarily responsible for the DNN-to-brain similarity. Similar to studies on pretrained models and brain similarity, AlKhamissi et al. (2024) investigated the reasons for the similarity of untrained language models and brain alignment by performing mechanistic interpretability of the models. By isolating components of the Transformer architecture (GPT-2 XL), they found that tokenization strategy and multihead attention are the two major components driving this better brain alignment.
@@ 6.4.4 Key Takeaways
      Contrary to previous claims in Shain et al. (2024), syntax and semantics are not associated with separate modules but instead share common brain language regions and are distributed across the language network.
@@ 6.5 Auditory Encoding
Figure 16: Brain prediction using self-supervised speech model: Data2Vec. The plot from Oota et al. (2023c) shows that speech-based models better predict early auditory cortex. Adapted with permission from the respective authors.
@@ 6.5.1 Alignment Between Pretrained Speech Models and Brains
Recently, several researchers have used popular deep learning models such as APC (Chung et al., 2020), Wav2Vec2.0 (Baevski et al., 2020), HuBERT (Hsu et al., 2021), and Data2Vec (Baevski et al., 2022) for encoding speech stimuli.  Millet et al. (2022) used a self-supervised learning model, Wav2Vec2.0, to learn latent representations of the speech waveform similar to human brain. They find that the functional hierarchy of its transformer layers aligns with the cortical hierarchy of speech in the brain and reveals the whole-brain organisation of speech processing with unprecedented clarity. This means that the first transformer layers map onto the low-level auditory cortices (A1 and A2), the deeper layers map onto brain regions associated with higher-level processes (e.g. STS and IFG). Vaidya et al. (2022) present the first systematic study to bridge the gap between recent four self-supervised speech representation methods (APC, Wav2Vec, Wav2Vec2.0, and HuBERT) and computational models of the human auditory system. Similar to Millet et al. (2022), they find that self-supervised speech models are the best models of auditory areas. Lower layers best modeled low-level areas, and upper-middle layers were most predictive of phonetic and semantic areas, while layer representations follow the accepted hierarchy of speech processing. Tuckute et al. (2023) analyzed 19 different speech models and found that some audio models derived in engineering contexts (model applications ranging from speech recognition and speech enhancement to audio captioning and audio source separation) produce poor predictions of auditory cortical responses, many task-optimized audio speech deep learning models outpredict a standard spectrotemporal model of the auditory cortex and exhibit hierarchical layer-region correspondence with auditory cortex. Further, Oota et al. (2023c) extended this analysis to more such deep learning based speech models (30 self-supervised speech models). They found that both language and auditory brain areas, are best aligned with intermediate layers in deep learning models. As shown in Figure 16, they also found that speech models better predict early auditory cortex than late language regions. Although pretrained speech models can understand broad aspects of speech in general, the implications of finetuning speech pretrained models for various speech-processing tasks for speech encoding in the brain, remain underexplored.
Figure 17: The pretrained Wav2Vec2.0 model and finetuned to eight different downstream speech tasks and their brain alignment. This figure is adapted from Oota et al. (2023d), and used with permission from the respective authors.
@@ 6.5.2 Underlying Speech Properties in Speech Models and Brains
Understanding the reasons behind the observed similarities between speech processing in speech models and brains can lead to more insights into both systems. Recent work Oota et al. (2023d) has found that using a finetuned Wav2Vec2.0 leads to improved brain alignment. In particular, as shown in Figure 17Oota et al. (2023d) build neural speech taskonomy models for brain encoding and aim to find speech-processing tasks that have the most explanatory capability of brain activation during naturalistic story listening experiments. They find that task-specific (Automated Speech Recognition (ASR), Entity Recognition (ER), Speaker Identification (SID) and Intent Classification (IC)) speech representations lead to a significant improvement in brain alignment compared to the pretrained Wav2Vec2.0 model for specific brain regions. Finetuning on ER, SID and IC leads to the best alignment for the early auditory cortex; finetuning on ASR provides the best encoding for the auditory associative cortex and language regions. Further, the layer-wise analysis of the effect of each speech task on the alignment with whole brain activity shows that the ASR task is better aligned in middle layers. Similar to language model fine-tuning with brain data in Schwartz et al. (2019)Li et al. (2024b) fine-tuned a pretrained Wav2Vec2.0 model with brain recordings and and found that this induced process modified the language representations, improving the model’s performance on downstream tasks from the SUPERB benchmark.
@@ 6.6.1 Alignment Between Vision Models and Brains
Similar to language, in vision, early models focused on independent models of visual processing (object classification) using CNNs (Yamins et al., 2014; Yamins & DiCarlo, 2016).  Eickenberg et al. (2017) use CNNs as candidate models to model human brain activity during the viewing of natural images by constructing predictive models based on their different CNN layers and BOLD fMRI activations. They find that there are similarities between the computations of convolutional networks and cognitive vision at the beginning and at the end of the ventral stream object-recognition process.  Cichy et al. (2016) further investigate the stages of human visual processing in both time (MEG recordings) and space (fMRI recordings). By comparing these findings with representations derived from deep neural networks (DNNs), the authors demonstrate that DNNs effectively encapsulate the sequential stages of human visual processing. This encompasses the progression from early visual areas towards the specialized pathways of the dorsal and ventral streams, highlighting the DNN’s capacity to mirror complex neural processes in both time and space. Despite the effectiveness of CNNs, it is difficult to draw specific inferences about neural information processing using CNN-derived representations from a generic object-classification CNN. Hence, Wang et al. (2019) built encoding models with individual feature spaces obtained from 21 computer vision tasks. One of the main findings is that features from 3D tasks, compared to those from 2D tasks, predict a distinct part of visual cortex. Recent efforts in visual encoding models, particularly self-supervised models (instance-prototype contrastive learning), operate by taking multiple samples over an image and projecting these through a deep convolutional neural network into a low-dimensional embeddings space (Konkle & Alvarez, 2022). The results show that these self-supervised models achieve parity with the category-supervised models in accounting for the structure of brain responses. Since the human visual system uses two parallel pathways for spatial processing and object recognition, while computer vision systems (CNNs) typically use a single pathway,  Choi et al. (2024) developed a dual-stream vision model to mimic human vision. This model uses two branches of CNNs to replicate the dorsal and ventral cortical pathways, aligning with the brain’s pathways and suggesting that distinct responses are driven more by visual attention and object recognition goals than by retinal input selectivity.
@@ 6.7 Multimodal Brain Encoding
Single modality stimulus. Here, participants engage in single modality stimuli, such as watching images or silent videos. Many brain encoding studies have focused on single modality stimuli, while representations are extracted from multimodal models (Oota et al., 2022e; Wang et al., 2023; Tang et al., 2024).  Oota et al. (2022e) experimented with multimodal models like CLIP, LXMERT, and VisualBERT and found VisualBERT better predict neural responses than vision-only models such as CNNs and Image Transformers. Similarly,  Wang et al. (2023) find that multimodal models like CLIP better predict neural responses in the visual cortex than previous vision-only models like CNNs. This is attributed to the fact that high-level human visual representations encompass semantics and the relational structure of the visual world beyond object identity (Gauthier et al., 2003). Recently, Tang et al. (2024) investigated a multimodal Transformer as the encoder architecture to extract the aligned concept representations for narrative stories and movies to model fMRI responses to naturalistic stories and movies, respectively. Since language and vision rely on similar concept representations, the authors perform a cross-modal experiment in which how well the language encoding models can predict movie-fMRI responses from narrative story features (story → movie) and how well the vision encoding models can predict narrative story-fMRI responses from movie features (movie → story). Overall, the authors find that cross-modality performance was higher for features extracted from multimodal transformers than for linearly aligned features extracted from unimodal transformers.
@@
Multimodality stimulus. Here, participants engage with multi-modal stimuli (e.g., watching movies that include audio). Recent studies have built encoding models where multi-modal stimulus representations are extracted using Transformer-based multi-modal models (Dong & Toneva, 2023; Nakagi et al., 2024).  Dong & Toneva (2023) present a systematic approach to probe multimodal video Transformer model by leveraging neuro-scientific evidence of multimodal information processing in the brain. The authors find that intermediate layers of a multimodal video transformer are better at predicting multimodal brain activity than other layers, indicating that the intermediate layers encode the most brain-related properties of the video stimuli. A recent study by Nakagi et al. (2024), which used fMRI during the viewing of 8.3 hours of video content, and discovered distinct brain regions associated with different semantic levels, highlighting the significance of modeling various levels of semantic content simultaneously. The video material was meticulously annotated in five distinct semantic categories—speech, object, story, summary, and time/place—employing advanced large language models to derive latent representations. These representations were then used to predict brain activity across the various semantic categories. The authors discovered that the lack of unique variance for Summary and TimePlace is a notable insight, suggesting that merely incorporating these types of information into encoding analyses may not adequately capture higher-level semantic representations in the brain.
@@ 7.3 Decoder Architectures
Figure 18: Scheme for Brain Decoding. Left: Image decoder (Smith, 2013), Right: Language Decoder (Wang et al., 2019).  The left Figure is adapted from (Smith, 2013) and the right Figure is adapted from (Wang et al., 2019). Adapted with permission from the respective authors.
@@ 7.4 Impact of Large Models on Brain Decoding
Masked autoencoders (He et al., 2022) play a vital role in learning low-rank representations by reconstructing missing parts of the input data. In the context of brain decoding, these models are often used for pretraining by masking out some brain voxels and attempting to reconstruct them, thereby learning the underlying representations (Chen et al., 2023; 2024c; Sun et al., 2024). These fMRI representations are then utilized as conditions for downstream conditional generation models, enhancing their ability to produce detailed and accurate reconstructions compared with linear models.
Contrastive learning (Khosla et al., 2020) has emerged as a powerful technique for representation learning by maximizing the similarity between related data points while minimizing the similarity between unrelated ones. This approach has been instrumental in aligning brain activity with corresponding stimuli in a shared embedding space, facilitating more accurate and semantically meaningful decoding. One of the most notable applications of contrastive learning in brain decoding is the CLIP model (Radford et al., 2021). CLIP aligns text and images in a shared embedding space, greatly enhancing the decoding of visual stimuli. These models decode brain activity into text descriptions that are then used to generate corresponding images, effectively bridging the gap between linguistic and visual representations. In brain decoding, researchers often align fMRI embeddings with CLIP-based embedding spaces, allowing for more precise and semantically rich reconstructions of visual stimuli from brain data (Chen et al., 2024c; Scotti et al., 2024).
@@
LLMs, particularly models in the GPT series (Brown et al., 2020), have revolutionized language decoding. These models are capable of generating coherent and contextually appropriate text based on brain activity patterns. For instance, instead of merely decoding vector representations of stimuli, recent studies have leveraged LLMs to reconstruct entire sentences or continuous language from fMRI data (Tang et al., 2023; Zhao et al., 2024). This shift from vector-based decoding to full text generation has significantly enhanced the semantic richness and contextual accuracy of the decoded output. The ability of LLMs to model complex language structures and generate text conditioned on neural data has opened new avenues for understanding how the brain processes language, providing practical applications in areas such as communication aids for individuals with speech impairments.
@@
Diffusion models (Ho et al., 2020), particularly those like Stable Diffusion (Rombach et al., 2022), have been pivotal in generating high-fidelity images from brain activity. These models leverage the noise-to-signal transformation process to produce detailed and semantically rich visual outputs. By conditioning these models on brain activity data, researchers have achieved remarkable success in reconstructing images that closely resemble the original stimuli (Scotti et al., 2024; Takagi & Nishimoto, 2022; Chen et al., 2023; 2024c; Ozcelik & VanRullen, 2023; Takagi & Nishimoto, 2023). The high resolution and fidelity of the generated images represent a significant improvement over previous methods, which often struggled to capture fine details and semantic accuracy simultaneously.
@@
Continuous Language, level2, l sep=10mm,[ (Affolter et al., 2020; Tang et al., 2023), tnode] ] ] [Auditory, level1, l sep=10mm, [Audio
@@
Reconstruction , level2, l sep=10mm,[ (Naselaris et al., 2009; Beliy et al., 2019; Takagi & Nishimoto, 2022; Ozcelik & VanRullen, 2023; Chen et al., 2023; Koide-Majima et al., 2024; Takagi & Nishimoto, 2023; Song et al., 2023; Benchetrit et al., 2024; Li et al., 2024a), tnode] ] [Video
Reconstruction, level2, l sep=10mm,[ (Nishimoto et al., 2011; Le et al., 2021; Kupershmidt et al., 2022; Chen et al., 2024c), tnode] ] ] ]
Figure 19: Comprehensive Categorization of Brain Decoding Studies Across Modalities and Tasks. This figure provides a systematic overview of brain decoding research, illustrating the diverse approaches used to reconstruct or classify stimuli from brain activity patterns. The categorization is organized by stimulus modalities (e.g., language, vision, audio, and multimodal) and specific tasks within each modality.
Authors Dataset Type Lang. Stimulus Representations ||||S|||| Dataset
@@
(Chen et al., 2024c) fMRI Latent Diffusion Model, CLIP 3 HCP fMRI-Video-Dataset
(Li et al., 2024a) EEG CLIP, Diffusion Model 10 Things EEG
(Song et al., 2023) EEG CLIP, ViT 10 Gifford2022 EEG
(Benchetrit et al., 2024) MEG CLIP 10 Things MEG